OpenStudy (anonymous):

if dy/dx = 2y^2 and if y=-1 when x=1, then when x=2, y=

OpenStudy (anonymous):

I know the answer if that helps.. -1/3 but I need to learn how to get there

OpenStudy (anonymous):

@Jhannybean help please

OpenStudy (anonymous):

getting anywhere?

OpenStudy (perl):

with wolfram I get http://www.wolframalpha.com/input/?i=solve+dy%2Fdx+%3D+2y^2%2C++y%281%29+%3D+-1 The particular solution is y = 1 / (1 - 2x) y(2) = 1 / ( 1 - 2*2 ) = 1 / (1 - 4) = 1 / -3 = -1/3

OpenStudy (perl):

@Jhannybean on the fourth line you introduced an x going over steps dy/(2y^2) = dx is correct the next step should be 1/2 * integral dy / y^2 = integral dx

OpenStudy (jhannybean):

Oh I see, I see.

OpenStudy (jhannybean):

Let me fix that.

OpenStudy (perl):

do you have to write all that over again? would be nice if we could just edit past posts

OpenStudy (perl):

or do you use a notepad file before you submit

OpenStudy (jhannybean):

the quote feature :)

OpenStudy (perl):

if its in Latex you can quote it ? i was wondering about that, i saw people quoting in red

OpenStudy (mendicant_bias):

Nvm, already covered.

OpenStudy (perl):

yep :)

OpenStudy (jhannybean):

Yeah, I arbitrarily added it in by mistake.

OpenStudy (perl):

no biggie. Errors, typos, errata, etc. are sometimes useful to learn from

OpenStudy (jhannybean):

\[\begin{align} \&:~\frac{dy}{2y^2} = dx \\&:~\int \frac{dy}{2y^2} = \int dx \\&:~\frac{1}{2}\int \frac{1}{y^2}dy = x + C \\&:~\frac{1}{2} \int y^{-2}dy =x+C \\&: ~ -\frac{1}{2y} =x+C \ \ \ \sf \text{when x = 1, y = -1} \\&: ~ \frac{1}{2} = 1 + C \\&: ~ \frac{1}{2}-1 = C \\&: ~ -\frac{1}{2} = C \\ \\&: ~ -\frac{1}{2y} = 2-\frac{1}{2} \\&: ~ -\frac{1}{2y} = \frac{3}{2} \\&:~ -2y =\frac{2}{3} \\&:~ \boxed{y=-\dfrac{1}{3}} \end{align}\]

OpenStudy (jhannybean):

There we go.

OpenStudy (perl):

^^